OpenStudy (anonymous):

Use implicit differentiation to find an equation of the tangent line to the curve at the given point. x^2/3 + y^2/3 = 4, (-3√3, 1)

OpenStudy (anonymous):

\[ \frac{2 x}{3}+\frac{2}{3} y y'=0 \]

OpenStudy (anonymous):

\[ y'=-\frac x y=-\frac{ -3\sqrt3}1= 3 \sqrt3 \]

OpenStudy (anonymous):

\[ y-1 = 3 \sqrt 3 (x + 3\sqrt 3)\\ y=3 \sqrt 3 x+ 28 \]